BobCat Inc. has cash of $22,000, accounts receivable of $78,000, inventory of $62,000, prepaid insurance 8,000 and land of $90,000. The company has only one current liability account- salary payable of $64,000. BobCat Inc. current ratio is 2.50?

True
False
Answer #1

Total current assets=Cash+AR+inventory+prepaid expenses

=(22,000+78,000+62,000+8000)=$170,000

Current ratio=Total current assets/ current liability

=$170,000/$64,000

=2.65625

Hence the statement is False.